# How Useful is Continued Pre-Training for Generative Unsupervised Domain Adaptation?

### v2

@@ 9 Conclusion
We introduce the setting of Generative UDA, and perform an investigation on the utility of continued pre-training in this setting. We compare the approach with the popular class of domain-invariance based methods for UDA, showing that CPT is both competitive with, and more stable than invariance based approaches. Our experiments show that the benefits of CPT extend to different architectures, tuning methods and data regimes. We motivate the need for target domain exposure through CPT by showing that performance on the target domain gradually degrades with increasing masking rate. Finally, we shed light on the interplay between masking and classification performance, and how this aids UDA. Our analysis shows that in predicting masked words that are informative to the downstream task, the model implicitly learns about the downstream task, furthering the benefits of directly learning the task. Our work connects the body of UDA research with that of instruction tuning, enabling an initial step towards a wider applicability of modern language models.
@@
In our study, we consider a class of PEFT methods that involve inserting learnable parameters between the layers of the model. Other classes of PEFT methods were not considered. However, we use Adapters and He et al. (2022) have shown connections between the method with Prefix Tuning  (Li & Liang, 2021) and LoRA (Hu et al., 2022).
@@ Appendix C Detailed results with the Amazon and MNLI Datasets
In addition to the MMD based method of Malik et al. (2023), we also compare CPT with other methods that promote domain invariance: 1) DANN (Ganin et al., 2016), which is the most widely used UDA method in NLP (Ramponi & Plank, 2020), but has been shown to be highly unstable; 2) CORAL (Sun et al., 2017), which minimizes second order statistics of the data embeddings. Additionally, with an emerging class of weight interpolation based methods, we make a comparison with task vector arithmetic Ilharco et al. (2022). The use of task vectors with PEFT methods beyond LoRA (Hu et al., 2022) has been unexplored in the literature, and we find that the method does not work with IA3. With fully fine-tuned models, the method improves in performance, but is still weaker than CPT.
@@ Appendix E PEFT Frameworks
are a popularly used and high performing PEFT framework, and He et al. (2022) have shown equivalence in the operations applied by Adapters, Prefix Tuning Li & Liang (2021) and LoRA Hu et al. (2022).
@@ Appendix I Single Phase CPT Training
The Maximum Mean Discrepancy (MMD) Gretton et al. (2012) measures the difference between first order moments of variables in a Reproducing Kernel Hilbert Space Aronszajn (1950). Multiple lines of work have shown that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, results in training instabilities and vanishing gradients Kashyap et al. (2021); Han & Eisenstein (2019).

### v3

@@ 9 Conclusion
We introduce the setting of Generative UDA, and perform an investigation on the utility of continued pre-training in this setting. We compare the approach with the popular class of domain-invariance based methods for UDA, showing that CPT is both competitive with, and more stable than invariance-based approaches. Our experiments show that the benefits of CPT extend to different architectures, tuning methods and data regimes. We motivate the need for target domain exposure through CPT by showing that performance on the target domain gradually degrades with increasing masking rate. Finally, we shed light on the interplay between masking and classification performance, and how this aids UDA. Our analysis shows that in predicting masked words that are informative to the downstream task, the model implicitly learns about the downstream task, furthering the benefits of directly learning the task. Our work connects the body of UDA research with that of instruction tuning, enabling an initial step towards a wider applicability of modern language models.
@@
In our study, we consider a class of PEFT methods that involve inserting learnable parameters between the layers of the model. Other classes of PEFT methods were not considered. However, we use Adapters and He et al. (2022) have shown connections between the method with Prefix Tuning  (Li and Liang, 2021) and LoRA (Hu et al., 2022).
@@ Appendix C Detailed results with the Amazon and MNLI Datasets
In addition to the MMD based method of Malik et al. (2023), we also compare CPT with other methods that promote domain invariance: 1) DANN (Ganin et al., 2016), which is the most widely used UDA method in NLP (Ramponi and Plank, 2020), but has been shown to be highly unstable; 2) CORAL (Sun et al., 2017), which minimizes second order statistics of the data embeddings. Additionally, with an emerging class of weight interpolation based methods, we make a comparison with task vector arithmetic Ilharco et al. (2022). The use of task vectors with PEFT methods beyond LoRA (Hu et al., 2022) has been unexplored in the literature, and we find that the method does not work with IA3. With fully fine-tuned models, the method improves in performance, but is still weaker than CPT.
@@ Appendix E PEFT Frameworks
are a popularly used and high performing PEFT framework, and He et al. (2022) have shown equivalence in the operations applied by Adapters, Prefix Tuning Li and Liang (2021) and LoRA Hu et al. (2022).
@@ Appendix I Single Phase CPT Training
The Maximum Mean Discrepancy (MMD) Gretton et al. (2012) measures the difference between first order moments of variables in a Reproducing Kernel Hilbert Space Aronszajn (1950). Multiple lines of work have shown that minimizing divergence measures like MMD, when combined with auxiliary task-specific loss functions, results in training instabilities and vanishing gradients Kashyap et al. (2021); Han and Eisenstein (2019).